Hakara Company has been using direct labor costs as the basis for assigning overhead to its many products. Under th allocation system, product A has been assigned overhead of $21.86 per unit, while product B has been assigned $13.35 per unit. Management feels that an ABC system will prov and has collected the following cost pool and cost driver information: ide a more accurate allocation of the overhead costs Ac tivity Driver Consumption 4,000 14,000 36, 000 Activity Costs Cost Drivers Cost Pools Machine setup Materials handling Electric power $284,000 Setup hours 70,000 Pounds of materials 72,000 Kilowatt-hours g cost information pertains to the production of A and B,just two of Hakaras many products: Number of units produced Direct materials cost Direct labor cost Number of setup hours Pounds of materials used Kilowatt-hours 5,000 10,000 $24,000 $34,000 $32,000 $35,000 200 2,000 2,000 2,000 4,000 100 Required 1. Use activity-based costing to determine a unit cost for each product. (Round your final answers to 2 decimal pl Cost per Unit Product A Product B
Answer #1

Answer:

Cost per unit

Product -A

15.42

Product -B

10.12

Working note for the above answer is as under:

Cost Pool

Activity cost$

Cost Driver

Driver consumption

Activity rate  

A

B

C=A/B

Machine set up

284000

Set up Hrs

4000

71

Materials handling

70,000

Pounds of material

14000

5

Electric Power

72,000

Kilowatt Hrs

36000

2

Overhead allocation to A &B

A

B

Cost Driver

Activity rate  

Driver cosumption

OH amount allocated

Driver cosumption

OH amount allocated

A

B

C=A*B

D

E=A*D

Set up Hrs

71

100

7100

200

14200

Pounds of material

5

2000

10000

2000

10000

Kilowatt Hrs

2

2000

4000

4000

8000

Total

21100

32200

Details

Product A

Product B

No Of units produced (A)

5,000

10,000

Cost Details

Direct Materials cost(B)

24000

34000

Direct Labor cost ©

32000

35000

Manufacturing overhead allocated   (D)

21100

32200

Total Cost(E=B+C+D)

77100

101200

Unit Cost =(E/A)

15.42

10.12